OpenStudy (anonymous):

What is the improper fraction to 12 2/41456? What is the improper fraction to 42 3/1242? What is the sum to those answers?

OpenStudy (texaschic101):

12 2/41456 = ((12 * 41456) + 2)/ 41456 = 497474/41456 42 3/1242 = ((42 * 1242) + 3)/ 1242 = 52167/1242 If I were you, to add these, I would turn them into decimals and then add them...rather then try finding common denominator. I had to cut this short because I lost track of time and I have to go. Sorry :(